In each reaction box, place the best reagent and cneed help with reagents and conditions!
Answer #1
Concepts and reason

The concept used to solve this problem is chemical reactions of dihalides and alkynes.

For a chemical reaction to occur, some substances are needed which are termed as reagents. These reagents react with the reactant molecules and initiate the chemical reaction in the reaction medium.

Fundamentals

Reagents are the substances which are used for the analysis or to bring a desire change in the chemical reaction.

Sodium amide is a strong base and it abstract the most acidic proton from the reactant molecule when added as a reagent in the chemical reaction. H2,Pt{{\rm{H}}_{\rm{2}}}{\rm{,Pt}} and sodium in liquid ammonia are used as reducing agents in the chemical reactions and water in acidic and in basic medium is used for oxidation of the compounds. mCPBA{\rm{mCPBA}} is a reagent which is used to carry out epoxidation of alkenes.

(1)

The reaction is as follows:

1s®
NaNH2(excess)

The reaction is as follows:

H2 , pt

The reaction is as follows:

N
140.490,cal) »
H2O,H2SO4(cat)
OH

The reagents involved in the given reaction sequence are shown as follows:

X
1) NaNH2(excess)
2) H2 , pt
3) H20,H2SO4(cat)
1
CI
OH

(2)

The reaction is as follows:

NaNH2, HC-CE
H3C—C=C-H
Na

The reaction is as follows:

O,
H3C—C=C Na
CH3CH2Br
H C—C=C—CH2CH3

The reaction is as follows:

Na, NH3 (1)
H3C-C=C—CH,CH;
CH2CH3
HC

The reaction is as follows:

CH12
СН3 ОН
ОН
Н
CH2CH3
KMnO4, NaOH,H,0,cold
н.
CH3 + НС
С-
Н4С
НО
ОН

The complete reaction is as follows:

СН3 ОН
ОН
H,CC=c—н —
1) NaNH,
2) CH,CHBr
—
3) Na,NH3 (1)
4) KMnO4, NaOH,H,O,cold
НА
CH3 + НС
HO
H
нон

(3)

The reaction is as follows:

H3C—C=C-H -
| NaNH,
H,C=c=cºNa

The reaction is as follows:

H3C—C=C Na
CH3CH2BrHzC—C=C—CH2CH3

The reaction is as follows:

Η
H3C—C=C—CH,CH, Hy/Lindars catalyst
CSC
CH,CH3

The reaction is as follows:

sullll
mCPBA
HCLCH, HzCV LCHz
HC
CH2CH3

The complete reaction is as follows:

1) NaNH2
2) CH3CH Br
H,C–C=C-H -
3) H/Lindars catalyst
4) mCPBA
+ HC
H3C
CH3

Ans: Part 1

The complete reaction is as follows:

X
1) NaNH2(excess)
2) H2 , pt
3) H20,H2SO4(cat)
1
CI
OH

Part 2

The complete reaction is as follows:

СН3 ОН
ОН
H,CC=c—н —
1) NaNH,
2) CH,CHBr
—
3) Na,NH3 (1)
4) KMnO4, NaOH,H,O,cold
НА
CH3 + НС
HO
H
нон

Part 3

The complete reaction is as follows:

1) NaNH2
2) CH3CH Br
H,C–C=C-H -
3) H/Lindars catalyst
4) mCPBA
+ HC
H3C
CH3